Abigail Lang and Colby Lohrey – Analyzing Water Runoff ...

Your presentation to the class was very good, and you had some good background information on the Blackstone River.

The final project paper is also direct and to the point. You did a good job with the measurements of the different categories of surfaces, and the discussion of you would define and compute the run-off index is also good. The main comment I have is that there are a number of other ways to set up reasonable methods there and some of them might be slightly better choices in some ways. For instance, while the boxplot would be one way to identify the “outlier” bad regions contributing the most to runoff, it might be even better with a large number of block measurements to do this: Take the 100 block RI values, compute the mean and SD of that data set. Then, after converting to equivalent z-scores, all regions having z-scores of 2 or higher (or 3 or higher) could be treated as “problem areas.” The reason this might be somewhat better in this case is that the z-score would reflect the actual runoff potential somewhat more faithfully than the “order statistics” from the boxplot, which deal more with the relative distribution of the RI values within the data set.

The other general comment I have is that the writing is still rather rough in spots. There are places where you overstate things and others where what you say is not too clear. See the specific comments below for some examples of each type.

Technical points in the citations and bibliography are handled reasonably well. You should be aware, though, that other professors here might not like to see such heavy reliance on Internet sources. Specific Comments

0) The whole opening section is rather overheated. I know you are trying to say that this runoff issue deserves attention because it is an important one. But there are ways to do that without bringing in almost every environmental problem you can think of.

1) Actually, I believe the main issue with regard to the Blackstone is that there is so much toxic material in the sediments on the river bed. A lot of this was deposited as a result of industrial uses of the river years ago. But the material is still there and is still causing a problem. The main issue with runoff water flowing in the river is that if that runoff is flowing quickly enough, then it can stir up those sediments and put them back into the water.

2) The sentence: “The impervious surfaces have being a major factor in water pollution due to the fact that it causes runoff water” (on page 2) is an example of what I meant by the “not too clear” comment above. It seems like you started to edit this but did not quite finish what you were trying to do. Also, impervious surfaces don’t “cause runoff water.” They increase the amount and speed of the water that runs off an area.

3) “contributing factors to water pollution that the human lifestyle in a community does” (page 3) does not really work as a phrase. I think you mean something like “Landscaping and lawncare, emissions from automobiles, construction, and septic systems, and general improper disposal of trash are all human contributions to water pollution.”

4) On page 3 again, the sentence “However a main issue of construction is more often than not, is eliminating pervious surfaces and replacing it with impervious surfaces.” is unclear and grammatically wrong. Saying it more directly willl help produce a clearer and sharper sentence: “Construction often eliminates pervious surfaces and replaces them with impervious ones.”

5) An arboretum is just an area containing a designated collection of trees. The Holy Cross campus is an arboretum, and the booklet you used as a source is, in effect, a catalog of those trees, together with other information about how the College manages and cares for them.

6) Individual action such as care in disposal of solid waste is a start, but efforts to manage a problem like the runoff problem at the College have to include more systemic changes too (like the ones you propose after that). So this was perhaps not the best place to start.

7) The idea about monitoring whose cars are leaking oil is an interesting one! Since there are no actual regulations about that in most localities, though, I can see that there might be some strong objections (“those darned tree-huggers are trying to tell me I can’t drive the car I want to drive!”) This would be another area where individual freedom comes into conflict with the needs of society!

Final Project Grades

Presentation: B+ (88)

Paper Content: B+ (88)

Paper Form: B+ (88)